# Frontiers in Intelligent Colonoscopy

### v1

@@ 4.4 VL models
$\bullet$ Remarks. Two possible reasons might explain the lag in VL research for colonoscopy. (a) Data-centric issue. The lack of well-structured and high-quality image-text pairs hinders progress. Future insights can be learnt from existing ideas. First, crawling unlabelled image-text data from social media [275] and scientific literature [276] can be used to build domain-specific foundation models. Second, language models such as GPT-4V [4] can generate diverse professional descriptions, offering an economical and scalable way to expand the knowledge space of the data. (b) Model-centric issue. Current VL techniques in colonoscopy have not kept pace, even with recent developments in multimodal language models (MLMs) [277, 8] for general domains. These techniques opt for a decoder-only strategy that unifies multiple tasks (e.g., detection, captioning) into a unified autoregressive framework (i.e., next-token prediction). These models are flexibly capable of processing input and output texts of varying lengths, without requiring additional task-specific prediction heads for different tasks.
5 Stepping into the Multimodal Land
@@
Recently, MLMs have demonstrated significant promise in leveraging language models to process multimodal signals, especially in “perceiving and interpreting” visual signals. Instruction tuning [278] is key to instructing the MLMs to execute domain-specific tasks aligned with user preferences. This section introduces three initiatives to advance multimodal research: how we create a large-scale instruction tuning dataset, ColonINST (§5.1) and how we train a colonoscopy-specific MLM, ColonGPT (§5.2). Finally, as in §5.3, we contribute a multimodal benchmark for conversational colonoscopy tasks, conduct diagnostic studies on ColonGPT, and share lessons from empirical observations.
5.1 Established instruction tuning dataset: ColonINST
@@
$\bullet$ Overview. Fig. (4-a) depicts our semi-automated workflow to create instruction tuning data in three steps. We begin by assembling colonoscopy images from public datasets. Following this, we use a category-specific prompt to interact with a multimodal AI chatbot, GPT-4V [4], yielding descriptive medical captions for these positive cases within the assembled data. Lastly, we reorganise the instruction tuning pairs derived from the data afore-prepared, enabling the model to perform four different colonoscopy tasks in an interactive, dialogue-based manner.
@@
$\bullet$ Data collection. As shown in Fig. (4-b), ColonINST contains 303,001 colonoscopy images, including 128,620 positive and 174,381 negative cases collected from 19 different data sources. To ensure data integrity and avoid label leakage, we establish a series of management rules to divide each dataset. For datasets with predefined divisions, such as KUMC [57], PICCOLO [39], WCE-CCDD [60], BKAI-Small [46], CP-CHILD [51], Kvasir-Instrument [50], and PS-NBI2K [16], we follow their original division rules. When such predefined rules are not available, we conform to widely recognised benchmarks, such as CVC-ClinicDB [21], CVC-ColonDB [19], ETIS-Larib [20], and the polyp category in Kvasir [28] according to the benchmark by Fan et al.[175], as well as positive samples in SUN-database [54] following the benchmark by Ji et al.[62]. For the remaining datasets (HyperKvasir [43], Kvasir-Capsule [56], CPC-Paired [47], Nerthus [29], GastroVision [66], EDD2020 [40], PolypGen [69], negative samples in SUN-database [54], remaining categories in Kvasir [28]), we allocate them proportionally as 60%/10%/30% for training/validation/testing purposes. With the above management rules, our final image division for is roughly 59.73%/8.77%/31.61%. As depicted in Fig. (4-c), all images are classified into a three-level structure, including 4 root/13 parent/62 child categories. In detail, the root level contains three positive categories: the pathological findings of various colonic diseases (110,970 cases); the anatomical structure related to the colon (5,511 cases); and therapeutic interventions related to colonoscopy (12,139 cases), covering the pre-operative, intra-operative, and post-operative stages. Targets outside our interest (e.g., stomach, oesophagus, normal Z-line, and gastric polyp not occurred during a colonoscopy) or lack of objects (e.g., normal mucosa, colon background) are classified under the negative root category (174,381 cases). We intentionally keep these negative samples, as they may be valuable for future gastrointestinal research. Fig. (4-d) presents a word cloud distribution for all categories names within ColonINST.
@@ 5.2 Proposed multimodal language model: ColonGPT
$\bullet$ Multigranularity multimodal adapter. Previous works [7, 8, 9] generally employ a multilayer perceptron architecture as a multimodal adapter, typically consisting of triple linear layers with intervening GELUs. However, handling all visual tokens introduces redundancy because not every token is equally significant, and it also incurs higher computational costs given the quadratic complexity in relation to the number of input tokens. To embrace these challenges, we propose a multimodal adapter that incorporates multigranularity pooling layers between two linear layers. As illustrated in the right of Fig. 6, we transform the embedding $\mathbf{E}_{v}\!\in\!\mathbb{R}^{\frac{HW}{P^{2}}\times d}$ from $d$-dim to $D$-dim using a linear layer followed by a GELU, then reshape it into the spatial format $\mathbf{F}_{v}\!\in\!\mathbb{R}^{\frac{H}{P}\times\frac{W}{P}\times D}$. To reduce the number of visual tokens while avoiding performance drops, we roll out three modifications for the pooling phase, each validated in §5.3.2. (a) Multigranularity views. We add a set of adaptive average pooling operations with $N$ kernel sizes $\{s_{1},\dots,s_{N}\}$ to obtain multigranularity features. In particular, this adaptive operator accommodates input sequences of varying lengths, i.e., the pooled feature for the kernel size $s_{n}$ is shaped as $\mathbb{R}^{s_{n}\times s_{n}\times D}$. (b) Positional encoding. Inspired by [281], we enhance the spatial information within each pooled feature by applying a 2D convolutional layer with the appropriate zero-padding setting. By default, a zero-pixel boundary is added around the input feature. (c) Global view. We also use a global average pooling layer with kernel size $1$ on the feature $\mathbf{F}_{v}$ to obtain a global view with the shape of $\mathbb{R}^{1\times 1\times D}$. Lastly, we reshape each pooled feature into flattened vectors: $\{\mathbf{F}_{p}^{n}\!\in\!\mathbb{R}^{s_{n}^{2}\times D}\}_{n=1}^{N}$ and $\mathbf{F}_{g}\!\in\!\mathbb{R}^{1\times D}$. We concatenate these vectors and process the resulting vector through the second linear layer to produce the final visual tokens $\mathbf{T}_{v}\!\in\!\mathbb{R}^{N_{v}\times D}$, where $N_{v}=(s_{1}^{2}+\cdots+s_{N}^{2}+1)$ denotes the length of the visual tokens.
$\bullet$ Implementation. Our model can be integrated with modern off-the-shelf visual encoders and LMs. To improve reproducibility for average users, we implement ColonGPT in an resource-friendly way. First, we employ SigLIP-SO (0.4B parameters) [5] as the visual encoder, with an input resolution of $H\!=\!W\!=\!384$, a patch size of $P\!=\!14$, and a visual embedding dimension of $d\!=\!1152$. This configuration yields a visual embedding $\mathbf{E}_{v}$ with a shape $\mathbb{R}^{729\times 1152}$, where the number of visual tokens is $729=\lfloor\frac{384}{14}\rfloor^{2}$. In addition, Phi-1.5 (1.3B parameters) [6] serves as the language tokenizer and language model, with a embedding size of $D\!=\!2048$. To reduce computational cost, the size of pooling kernels is set to $\{s_{1},s_{2}\}\!=\!\{14,7\}$, significantly reducing the visual tokens $N_{v}$ from $729$ to $246$, a reduction of $66.26\%$. This design allows us to complete training in five hours, facilitating rapid proof-of-concept development.
$\bullet$ Training recipe. We implement our model using PyTorch library, accelerated by four NVIDIA A100-40GB GPUs. The AdamW optimiser is used with an initial learning rate of 2e-3 and a cosine learning rate scheduler. Our ColonGPT is trained on the combination of training and validation dialogues from ColonINST. The complete training runs for three epochs, with a batch size of 16 per GPU and a gradient accumulation every 2 steps. During training, the visual encoder is frozen, focusing on the trainable multimodal adapter and the LM. For efficiency, we utilise the low-rank adaptation (LoRA [282]) strategy for LM, with a rank of $r\!=\!128$ and a scaling factor of $\alpha\!=\!256$.
TABLE VII: Multimodal benchmark for three conversational tasks. “LoRA” refers to fine-tuning using low-rank adaptation [282]. “EXT” indicates the use of pre-trained weights on extra data. We compare the results on the seen samples from the validation set and the unseen samples from the testing set of ColonINST. The symbol $\uparrow$ signifies that a higher score reflects better performance.
@@
ColonGPT (Ours) SigLIP-SO (384px/link) Phi1.5 (1.3B/link) - 94.02% 85.81% 99.02% 83.42% 65.89% 45.77%
@@
5.3 Experiments
5.3.1 Multimodal benchmark
$\bullet$ Model competitors. To establish a widely accepted multimodal benchmark for the community, we select eight popular MLMs as competitors, including six general-purpose and two medically tailored models. As shown in Tab. VII, each competitor has two training setups depending on whether it uses LoRA [282] or initialises knowledge from additional pre-training data. We retrain all competitors using the combined training and validation dialogues from ColonINST, as used in our ColonGPT.
$\bullet$ Evaluation protocols. We quantitatively evaluate three conversational tasks for the multimodal benchmark. For the two classification-based tasks, namely CLS and REG, we adopt the accuracy metric ($\mathcal{A}$) to calculate the ratio of correctly predicted categories to the total number of predictions. For the REC task, we use the intersection over union (IoU) metric to measure the localisation precision. Furthermore, due to the subjective nature of language in the CAP task, we qualitatively analyse the medical accuracy of the responses by verifying the correct identification of the anatomical structures and category names visible in the images, or relevant clinical information descriptions.
$\bullet$ Learning ability. We begin by conducting an open-book test for each model to quantitatively measure how effectively each model has internalised the visual and linguistic patterns from the training phase. Specifically, we evaluate each model on the samples they have seen during training, i.e., 27,606 validation dialogues. The “seen” columns in Tab. VII show that we achieve the highest scores in the CLS ($\mathcal{A}\!=\!94.02\%$) and REC ($\text{IoU}\!=\!65.89\%$) tasks. This implies that ColonGPT has a better learning ability, which allows it to classify images and understand reference expressions related to specific visual regions. Furthermore, we achieve an accuracy of $99.02\%$, below $99.32\%$ achieved by LLaVA-v1.5 (#C2). This gap ($\Delta\!=\!0.3\%$) is due to a 7B-level LM used in LLaVA-v1.5, which has also been pretrained on additional data and thus gained more knowledge.
$\bullet$ Generalisation ability. We further conduct a closed-book test to examine each model’s ability to generalise knowledge to unseen conditions, i.e., 165,830 testing samples of ColonINST. The “unseen” columns in Tab. VII consistently reveal our superior performance in unseen samples in all three tasks. Recall that our model performs worse than LLaVA-v1.5 in the REG task for seen scenarios. However, this gap is overtaken when exposed to unseen scenarios, where we achieve an accuracy of $83.42\%$, even exceeding the 7B-level LLaVA-v1.5 (#C2) by $10.54\%$. Compared to the runner-up model, LLaVA-Med-v1.5 (#H1), our model surpasses it by $1.2\%$ in the seen scenarios for the REC task, and this gap is further widened by $3.8\%$ in the unseen scenarios. These results show the potential of ColonGPT as a conversational assistant for multimodal colonoscopy tasks, especially in generalising to unseen data.
@@ 5.3.1 Multimodal benchmark
ConvNeXtV2-L 384px/link 82.95% 78.63% 33.74%
ViT-L 384px/link 82.16% 77.04% 40.78%
MAE-L* 384px/link 80.85% 75.87% 38.53%
MAE-L 224px/link 81.95% 77.62% 43.25%
DINOv2-L* 384px/link 35.03% 22.91% 6.79%
DINOv2-L 224px/link 21.22% 7.96% 2.69%
CLIP-L 336px/link 83.99% 78.67% 41.54%
SigLIP-SO 384px/link 85.81% 83.42% 45.77%
@@
full-tuning - - 78.06% 73.79% 50.20%
LoRA 4 8 85.43% 82.75% 45.02%
LoRA 8 16 84.45% 80.78% 44.98%
LoRA 16 32 84.39% 80.81% 45.90%
LoRA 32 64 84.91% 82.73% 45.56%
LoRA 64 128 83.84% 81.19% 43.57%
LoRA 128 256 85.81% 83.42% 45.77%
LoRA 256 512 82.93% 79.96% 48.27%
@@
$\bullet$ Qualitative analysis. Fig. 7 illustrates our model’s three multimodal abilities across four conversational tasks. (a) Comprehension ability: In the CLS task, we identify subtle visual features, distinguishing “high grade dysplasia” from “adenoma” in visually similar images. In the REG task, we correctly translate complex visual features from the given coordinates into precise medical terminology. (b) Localisation ability: This entails ColonGPT understanding language query and localising visual target within a complex colon environment. The outputs of the REC task showcase ColonGPT’s precision in localising specified expressions using bounding boxes. (c) Captioning ability: This requires the synthesis of visual information into coherent, clinically relevant text. Our model provides descriptions of a pedunculated polyp, detailing its morphology, contextual characteristics, and potential clinical relevance. Additionally, ColonGPT can describe the treatment procedure when an instrument is present, e.g., “The polyp is being examined using a snare.”
5.3.2 Diagnostic studies

### v2

@@ IV-D VL models
$\bullet$ Remarks. Two possible reasons might explain the lag in VL research for colonoscopy. (a) Data-centric issue. The lack of well-structured and high-quality image-text pairs hinders progress. Future insights can be learnt from existing ideas. First, crawling unlabelled image-text data from social media [275] or the scientific literature [276] can be used to build domain-specific foundation models. Second, language models such as GPT-4V [4] can generate diverse professional descriptions, offering an economical and scalable way to expand the knowledge space of the data. (b) Model-centric issue. Current VL techniques in colonoscopy have not kept pace, even with recent developments in multimodal language models (MLMs) [277, 8] for general domains. These techniques opt for a decoder-only strategy that unifies multiple tasks (e.g., detection, captioning) into a unified autoregressive framework (i.e., next-token prediction). These models are flexibly capable of processing input and output texts of varying lengths, without requiring additional task-specific heads for different tasks.
V Advancing Multimodal Learning for Colonoscopy
@@
Recently, MLMs have demonstrated significant promise in leveraging language models to process multimodal signals, especially in “perceiving and interpreting” visual signals. Instruction tuning [278] is key to instructing the MLMs to execute domain-specific tasks aligned with user preferences. This section introduces three initiatives to advance multimodal research: how we create a large-scale instruction tuning dataset, ColonINST (§V-A) and how we train a colonoscopy-specific MLM, ColonGPT (§V-B). Finally, we contribute the first multimodal benchmark for conversational tasks (§V-C), conduct diagnostic studies on ColonGPT (§V-D), and share lessons from our empirical observations (§V-E).
V-A Established instruction tuning dataset: ColonINST
@@
Fig. (4-a) depicts our semi-automated workflow to create instruction tuning data in three steps. We begin by assembling colonoscopy images from public datasets. Following this, we use a category-specific prompt to interact with a multimodal AI chatbot, GPT-4V [4], yielding descriptive medical captions for these positive cases within the assembled data. Lastly, we reorganise the instruction tuning pairs derived from the data afore-prepared, enabling the model to perform four different colonoscopy tasks in an interactive, dialogue-based manner.
@@
$\bullet$ Data collection. As shown in Fig. (4-b), ColonINST contains 303,001 colonoscopy images, including 128,620 positive and 174,381 negative cases collected from 19 different data sources. To ensure data integrity and avoid label leakage, we establish a series of management rules to divide each dataset. For datasets with predefined divisions, such as KUMC [57], PICCOLO [39], WCE-CCDD [60], BKAI-Small [46], CP-CHILD [51], Kvasir-Instrument [50], and PS-NBI2K [16], we follow their original division rules. When such predefined rules are not available, we conform to widely recognised benchmarks, such as CVC-ClinicDB [21], CVC-ColonDB [19], ETIS-Larib [20], and the polyp category in Kvasir [28] according to the benchmark by Fan et al.[175], as well as positive samples in SUN-database [54] following the benchmark by Ji et al.[62]. For the remaining datasets (HyperKvasir [43], Kvasir-Capsule [56], CPC-Paired [47], Nerthus [29], GastroVision [66], EDD2020 [40], PolypGen [69], negative samples in SUN-database [54], remaining categories in Kvasir [28]), we allocate them proportionally as 60%/10%/30% for training/validation/testing purposes. With the above management rules, our final image division for is roughly 59.73%/8.77%/31.61%. As depicted in Fig. (4-c), all images are classified into a three-level structure, including 4 root/13 parent/62 child categories. In detail, the root level contains three positive categories: the pathological findings of various colonic diseases (110,970 cases); the anatomical structure related to the colon (5,511 cases); and therapeutic interventions related to colonoscopy (12,139 cases), covering the pre-operative, intra-operative, and post-operative stages. Targets outside our interest (e.g., stomach, oesophagus, normal Z-line, and gastric polyp not occurred during a colonoscopy) or lack of objects (e.g., normal mucosa, colon background) are classified under the negative root category (174,381 cases). We intentionally keep these negative samples, as they may be valuable for future gastrointestinal research. As shown in Fig. (4-d), we present a word cloud distribution for all categories names within ColonINST.
@@ V-B Proposed multimodal language model: ColonGPT
$\bullet$ Multigranularity multimodal adapter. Previous works [7, 8, 9] generally employ a multilayer perceptron architecture as a multimodal adapter, typically consisting of triple linear layers with intervening GELUs. However, handling all visual tokens introduces redundancy because not every token is equally significant, and it also incurs higher computational costs given the quadratic complexity in relation to the number of input tokens. To embrace these challenges, we propose a multimodal adapter that incorporates multigranularity pooling layers between two linear layers. As illustrated in the right of Fig. 6, we transform the embedding $\mathbf{E}_{v}\!\in\!\mathbb{R}^{\frac{HW}{P^{2}}\times d}$ from $d$-dim to $D$-dim using a linear layer followed by a GELU, then reshape it into the spatial format $\mathbf{F}_{v}\!\in\!\mathbb{R}^{\frac{H}{P}\times\frac{W}{P}\times D}$. To reduce the number of visual tokens while avoiding performance drops, we roll out three modifications for the pooling phase, each validated in §V-D. (a) Multigranularity views. We add a set of adaptive average pooling operations with $N$ kernel sizes $\{s_{1},\dots,s_{N}\}$ to obtain multigranularity features. In particular, this adaptive operator accommodates input sequences of varying lengths, i.e., the pooled feature for the kernel size $s_{n}$ is shaped as $\mathbb{R}^{s_{n}\times s_{n}\times D}$. (b) Positional encoding. Inspired by [281], we enhance the spatial information within each pooled feature by applying a 2D convolutional layer with the appropriate zero-padding setting. By default, a zero-pixel boundary is added around the input feature. (c) Global view. We also use a global average pooling layer with kernel size $1$ on the feature $\mathbf{F}_{v}$ to obtain a global view with the shape of $\mathbb{R}^{1\times 1\times D}$. Lastly, we reshape each pooled feature into flattened vectors: $\{\mathbf{F}_{p}^{n}\!\in\!\mathbb{R}^{s_{n}^{2}\times D}\}_{n=1}^{N}$ and $\mathbf{F}_{g}\!\in\!\mathbb{R}^{1\times D}$. We concatenate these vectors and process the resulting vector through the second linear layer to produce the final visual tokens $\mathbf{T}_{v}\!\in\!\mathbb{R}^{N_{v}\times D}$, where $N_{v}=(s_{1}^{2}+\cdots+s_{N}^{2}+1)$ is the length of visual tokens.
$\bullet$ Implementation. Our model can be integrated with modern off-the-shelf visual encoders and LMs. To improve reproducibility for average users, we implement ColonGPT in a resource-friendly way. First, we employ SigLIP-SO (0.4B parameters) [5] as the visual encoder, with an input resolution of $H\!=\!W\!=\!384$, a patch size of $P\!=\!14$, and a visual embedding dimension of $d\!=\!1152$. This configuration yields a visual embedding $\mathbf{E}_{v}$ with a shape $\mathbb{R}^{729\times 1152}$, where the number of visual tokens is $729=\lfloor\frac{384}{14}\rfloor^{2}$. In addition, Phi-1.5 (1.3B parameters) [6] serves as the language tokenizer and language model, with a embedding size of $D\!=\!2048$. To reduce computational cost, the size of pooling kernels is set to $\{s_{1},s_{2}\}\!=\!\{14,7\}$, significantly reducing the visual tokens $N_{v}$ from $729$ to $246$, a reduction of $66.26\%$. This design allows us to complete the training in seven hours, facilitating rapid proof-of-concept development.
@@
TABLE VII: Multimodal benchmark for three conversational tasks. “LoRA” refers to fine-tuning using low-rank adaptation [282]. “EXT” indicates the use of pre-trained weights on extra general data. We compare the results on the seen samples from the validation set and the unseen samples from the testing set of ColonINST. Symbol $\uparrow$ means that a higher score reflects better performance.
@@
ColonGPT (Ours) SigLIP-SO (384px/link) Phi1.5 (1.3B/link) - 94.06% 83.24% 99.96% 80.18% 85.74% 56.24%
$\bullet$ Training recipe. We implement our model using the PyTorch library, accelerated by two NVIDIA H200 GPUs. The AdamW optimiser is used with an initial learning rate of 2e-3 and a cosine learning rate scheduler. Our reproducible process consists of two training stages. (a) Pre-alignment stage: First, we train the multigranularity multimodal adapter on $\sim$83K image-caption pairs to pre-align the visual encoder with the language model, enabling it to interpret visual tokens. At this stage, both the visual encoder and the LM remain frozen, while the focus is on the multimodal adapter with a learning rate of 2e-4. (b) Supervised fine-tuning stage: Next, we adapt the model to address colonoscopy-related tasks by fine-tuning with CLS, REG, and REC data, amounting to $\sim$202K image-text pairs. For efficiency, the low-rank adaptation strategy (LoRA [282]) is applied to the LM, with a rank of $r\!=\!128$ and a scaling factor of $\alpha\!=\!256$. During this stage, we train the adapter and LM with a learning rate of 2e-3 and 2e-4, respectively. Each stage lasts three epochs with a batch size of 16 per GPU and gradient accumulation every two steps. The duration of training takes about 1.4 hours for pre-alignment and 5.3 hours for fine-tuning.
V-C Multimodal benchmark
$\bullet$ Model competitors. To establish a widely accepted multimodal benchmark for the community, we select eight popular MLMs as competitors, including six general-purpose and two medically tailored models. As shown in Tab. VII, each competitor has two training setups depending on whether it uses LoRA [282] or initialises knowledge from weights pre-trained on additional general data. We retrain each competitor using all the training and validation dialogues from ColonINST.
$\bullet$ Evaluation protocols. We quantitatively evaluate three conversational tasks for the multimodal benchmark. For the two classification-based tasks, namely CLS and REG, we adopt the accuracy metric ($\mathcal{A}$) to calculate the ratio of correctly predicted categories to the total number of predictions. For the REC task, we use the intersection over union (IoU) metric to measure the localisation precision. Furthermore, due to the subjective nature of language in the CAP task, we qualitatively analyse the medical accuracy of the responses by verifying the correct identification of the anatomical structures and category names visible in the images, or relevant clinical descriptions.
$\bullet$ Learning ability. We begin by conducting an open-book test for each model to quantitatively measure how effectively each model has internalised the visual and linguistic patterns from the training phase. Specifically, we evaluate each model on the samples they have seen during training, i.e., validation dialogues in ColonINST. The “seen” columns in Tab. VII show that we achieve the highest scores in the CLS ($\mathcal{A}\!=\!94.06\%$), REG ($\mathcal{A}\!=\!99.96\%$), and REC ($\text{IoU}\!=\!85.74\%$) tasks. This suggests that our ColonGPT has better learning ability, which allows it to correctly classify images and understand reference expressions related to specific visual regions.
$\bullet$ Generalisation ability. We further conduct a closed-book test to examine each model’s ability to generalise knowledge to unseen conditions, i.e., testing samples of ColonINST. The “unseen” columns in Tab. VII consistently reveal our superior performance on unseen samples across all three tasks. Recall that our model slightly outperforms LLaVA-v1.5 in the REG task for the seen scenarios, i.e., $99.96\%$ (Ours) vs. $99.32\%$ (#C2). However, this gap is further widened when exposed to unseen scenarios, where we achieve an accuracy of $80.18\%$, even exceeding the 7B-level LLaVA-v1.5 (#C2) by $7.3\%$. Moreover, compared to medically tailored models, ColonGPT shows great potential as a colonoscopy-specific conversational assistant for multimodal tasks, especially in generalising to unseen data. For example, we exceed #G1 and #H1 models by $5.2\%$ and $4.0\%$, respectively in the unseen scenarios for the CLS task.
@@ V-C Multimodal benchmark
ConvNeXtV2-L 384px/link 81.51% 76.41% 53.09%
ViT-L 384px/link 83.99% 78.79% 56.64%
MAE-L* 384px/link 83.30% 79.51% 53.59%
MAE-L 224px/link 82.85% 79.19% 57.57%
DINOv2-L 224px/link 22.48% 10.09% 5.65%
CLIP-L 336px/link 77.05% 71.92% 51.68%
SigLIP-SO 384px/link 83.24% 80.18% 56.24%
@@
full-tuning - - 70.03% 63.07% 13.17%
LoRA 8 16 80.51% 77.61% 51.61%
LoRA 16 32 81.95% 77.87% 52.41%
LoRA 32 64 82.63% 78.76% 52.11%
LoRA 64 128 82.27% 79.72% 52.20%
LoRA 128 256 83.24% 80.18% 56.24%
LoRA 256 512 81.75% 77.89% 54.03%
LoRA 512 1024 79.39% 74.54% 55.52%
@@
$\bullet$ Qualitative analysis. Fig. 7 illustrates our model’s three multimodal abilities across four conversational tasks. (a) Comprehension ability: In the CLS task, we identify subtle visual features, distinguishing “high grade dysplasia” from “adenoma” in visually similar images. In the REG task, we correctly translate complex visual features from the given coordinates into precise medical terminology. (b) Localisation ability: This entails ColonGPT understanding language query and localising visual target within a complex colon environment. The outputs of the REC task showcase ColonGPT’s precision in localising specified expressions using bounding boxes. (c) Captioning ability: This requires the synthesis of visual information into coherent, clinically relevant text. Our model provides descriptions of a pedunculated polyp, detailing its morphology, contextual characteristics, and potential clinical relevance. Additionally, ColonGPT can describe the treatment procedure when an instrument is present, e.g., “The presence of a snare, a tool used to remove polyp”
V-D Diagnostic experiments